OpenStudy (anonymous):

f(x)=3 sqaure root x+3

OpenStudy (anonymous):

\[f(x)=3\sqrt{x+3}\]

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

\[f(x)=3\sqrt{x+3}\]

OpenStudy (anonymous):

Do you want to solve for x?

OpenStudy (anonymous):

find an equation for inverse function

OpenStudy (anonymous):

Rewrite as \[x = 3\sqrt{f(x)+3}\] Now solve for f(x)

OpenStudy (anonymous):

Your result will be the inverse.

OpenStudy (anonymous):

Yep, as said above ^^ sorry for the long wait, I didn't realise you had responded!

OpenStudy (anonymous):

Sara no offense but you're kinda cute :) haha

OpenStudy (anonymous):

haha thanks >.< not really sure how I could take offense at that though! @dhrumil91 do you understand that?

OpenStudy (anonymous):

Skype: ;) DeutschArzt Your choice!